Prior to liquidating their partnership, Ellis and Montgomery had capital accounts of $71,000 and $120,000, respectively. Prior to liquidation, the partnership had no cash assets other than what was realized from the sale of assets. These partnership assets were sold for $167,000. The partnership had $6,000 of liabilities. Ellis and Montgomery share income and losses equally. Determine the amount received by Ellis as a final distribution from liquidation of the partnership.

$
Answer #1

Capital, Ellis = $71,000

Capital, Montgomery = $120,000

Liabilities = $6,000

Sale price of assets = $167,000

Loss on disposal of assets = (Capital, Ellis + Capital, Montgomery + Liabilities) - Sale price of assets

= (71,000 + 120,000 + 6,000) - 167,000

= 197,000 - 167,000

= $30,000

Loss to be shared by Ellis = 30,000 x 1/2

= $15,000

Amount received by Ellis as a final distribution from liquidation of the partnership =  Capital, Ellis - Loss to be shared by Ellis

= 71,000 - 15,000

= $56,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.